you are given 4 vectors with equal magnitudes. under what conditions will their sum be zero? if a vector has an equal, positive x and y components, what are the angles it makes with the axes?

A superhero jumped horizontally from a 31 m high rooftop to the ground immediately after he spotted a child walking towards a busy street all alone. Fortunately, he lands on the street just next to the child in 2.5s and the child is rescued! If his initial speed is 22 m/s. how far the child was encountered from the wall of the building as the superhero caught him? What was his speed of impact to the ground?

A motorist traveling at a constant speed of 15m/s passes a school crossing corner where the speed limit is 10m/s. a police officer on a motorcycle, waiting at the corner, starts off in pursuit with constant acceleration of 3m/s2. how long should it take for the officer to catch up the motorist? what is the distance traveled by the officer while in pursuit?

you throw a ball upward from rooftop with a velocity of 15m/s. If the ball finally hits the ground with a downward velocity of 28m/s, how tall the building is?

which of the following objects is not accelerating? 1) a jogger moving with a constant speed 2) bicyclist moving with a constant speed southbound 3) a car making a u-turn with a constant speed.

An automobile, traveling at a constant speed of 25m/s, enters 90 degree curve and emerges from this curve 6s later. what are the x and y components of average acceleration during this time interval?

While jogging this morning, you encountered a circular track. as you finished one round along the track with uniform speed, what were your average velocity and average acceleration for this motion?

A motorcycle is traveling clockwise around a traffic circle at a steady speed of 50km/h. The x axis is considered eastward and y axis is northward. What are the x and y components of velocity at a point when the motorcycle is at the eastern point of circle?

A tennis ball initially traveling horizontally with velocity of 13m/s, collides with a wall and bounces back with the same speed. If the collision lasts for 0.4s, what is the magnitude and direction of the acceleration during the collision?

a champion discus thrower throws a discus with an initial speed of 26m/s. if the discus is thrown upward at and angle of 45 degree, what maximum height and what range will the discus attain? what is it's time of flight? Ignore the height of the hand over the ground.

Derive the x- and y-co-ordinates of a projectile at a time T after it has been projected at an angle theta with the horizon, with an initial speed V.
Answer #1

DO NOT POST THESE MANY QUESTIONS AT ONCE............FOLLOW THE RULES, PLEASE

The only requirement for vectors to add up to a zero vector is that the sums of all the components in the same direction of each other is zero

If vector has equal x and y components, then angle will be 45 degree

--------------------------------------------------------------------------------------------------

x = vt

x = 22 * 2.5 = 55 m from wall of building

vy = sqrt (2gh)

vy = 24.65 m/s

so,

impact speed, v = sqrt (24.652 + 222) = 33.04 m/s

-----------------------------------------------------------------------------

equating distance covered by police officer and distance covered by motorist

1/2at2 = vt

t = 2v / t

t = 2 * 15 / 3

t = 10 seconds

distance traveled in pursuit by officer = 150 m

--------------------------------------------------------------------